A 2.8 kg block slides on a frictionless 20 degree inclined plane. A force of 23 N acting parallel to the incline and up the incline is applied to the block. The block is moving upward. What is the acceleration of the block?


1
Expert's answer
2021-06-18T12:33:21-0400

Let's apply Newton's Second Law of Motion:


"F-mgsin\theta=ma,""a=\dfrac{F-mgsin\theta}{m},""a=\dfrac{23\ N-2.8\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot sin20^{\circ}}{2.8\ kg}=4.86\ \dfrac{m}{s^2}."